OpenStudy (anonymous):

4/5+1/7= what is is asking to do is estimate sums and differences of fractions and mixed numbers. The example is when u add or subtract fractions, ask: Is each fraction closest to 0,to 1/2 or to 1 can u show me in steps and ex plane please.

OpenStudy (anonymous):

4/5 = 8/10 = .8, around 1. 1/7 = .142 = around 0 1 + 0 = 1 The actual sum is .94, so our guess of 1 is good.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how did u get a decimal for 1/7

OpenStudy (anonymous):

for it to round to 0